Lea’s school is selling tickets for the upcoming choral performance. On The first day of ticket sales, the school sold 10 adult tickets and 10 child tickets for a total of $200. On the second day, the school sold 13 adult tickets And 2 child tickets for a total of $172. How much did each child ticket cost?
Found 2 solutions by josgarithmetic, greenestamps:
Answer by josgarithmetic(39617)

simple linear system two-variable application

Answer by greenestamps(13200)



With basic algebra....

10a+10c = 200
13a+2c = 172

Solve by any method you like. Note that the first equation is equivalent to a+c = 20; whatever solution method you use, that will make things easier.

-------------------------------------------------------------------

The same thing, doing mostly the same calculations as with algebra, but informally, using logical reasoning instead.

(1) 10 adults plus 10 children equals $200, so 2 adults plus 2 children equals $200/5 = $40.
(2) Compare that to 13 adults plus 2 children equals $172; the difference is 11 adults for $132.
So each adult ticket costs $132/11 = $12; then each child ticket costs $8.